# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ 3.3.1 Robust Confidence Sets
The following theorem ensures that under the null hypothesis, $\hat{Q}(\beta(a),a)$ is asymptotically normal.
@@
Under Assumptions 2, 3, 4, 5 and 6, we have that under $H_{0}:\beta(a)=\beta_{0}(a)$
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(a),a)}{\sigma_{Q}(a)}\to\mathcal{N}(0,1)\text{ in distribution}.$
Moreover, the asymptotic variance of $\hat{Q}(\beta(a),a)$ can be consistently estimated, and hence, the approach for robust confidence sets is valid.
@@
Under the assumptions of Theorem 12 we have that under $H_{0}:\beta(a)=\beta_{0}(a)$
$\frac{\widehat{SE}_{Q}^{2}(\beta_{0}(a),a)}{\sigma_{Q}^{2}(a)}\to 1\text{ in probability.}$.
@@
For the validity of the robust confidence sets, one still needs Assumption 5 which does not allow for IV that are too weak (see also Remark 3). However, we do not need the additional conditions $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(a),\,\|\phi-\hat{\phi}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(a)=o_{P}(h^{1/2})$ that – given a rate of convergence for $\hat{f}^{k}$ and $\hat{\phi}^{k}$ – pose even stronger restrictions on the IV-strength $\sigma_{Q}^{2}(a)$. In that sense, the robust confidence set construction works under slightly less restrictive conditions than the asymptotic normality of the point estimator.
@@ 4 Applications
We evaluate our method on simulated and real-world datasets. The method is implemented in the R package IVDML.111available from https://github.com/cyrillsch/IVDML The code to reproduce the simulation and real data results is available on GitHub. [^2] https://github.com/cyrillsch/IVDML_Application Before considering simulations and real data applications, we point out an aspect that is important when applying the methods in practice.
@@ 4.1 Repeated Cross-Fitting
The partition of the dataset into the $K$ folds is random. The finite sample estimators $\hat{\beta}$ and $\hat{\sigma}^{2}$ for the homogeneous treatment effect and $\hat{\beta}(a)$ and $\hat{\sigma}^{2}(a)$ for the heterogeneous treatment effect depend on this random partition. Let us focus on the homogeneous treatment effect in the following (the aggregation for the heterogeneous treatment effect is analogous). Following [12], the procedure can be repeated $S$ times with independent random partitions, yielding estimators $(\beta_{s})_{s=1,\ldots,S}$ and $(\hat{\sigma}^{2}_{s})_{s=1,\ldots,S}$. The point estimator is then aggregated using the median, $\hat{\beta}^{*}=\mathrm{median}\{\hat{\beta}_{s};\,s=1,\ldots,S\}$ and the asymptotic variance additionally is updated by a term incorporating the sample variability, $\hat{\sigma}^{2,*}=\mathrm{median}\{\hat{\sigma}_{s}^{2}+(\hat{\theta}_{s}-\hat{\theta}^{*})^{2};\,s=1,\ldots,S\}$. For the robust confidence set construction, we use the same idea. Let $\left(\hat{Q}_{s}(\cdot)\right)_{s=1,\ldots,S}$ and $\left(\widehat{SE}_{Q,s}^{2}(\cdot)\right)_{b=1,\ldots,B}$ be obtained from (13) and (14). Define $\hat{Q}^{*}(\cdot)=\mathrm{median}\left\{\hat{Q}_{s}(\cdot);\,s=1,\ldots,S\right\}$ and $\widehat{SE}_{Q}^{2,s}(\cdot)=\mathrm{median}\left\{\widehat{SE}_{Q,s}^{2}(\cdot)+\left(\hat{Q}_{s}(\cdot)-\hat{Q}^{*}(\cdot)\right)^{2};\,s=1,\ldots,S\right\}$, where the medians are taken argument-wise.
@@ 4.2 Simulations
$Z_{i}$ $\leftarrow 0.5X_{i}+E_{Z,i}$
$\delta_{i}$ $\leftarrow 0.7H_{i}+0.7E_{\delta,i},$
$\epsilon_{i}$ $\leftarrow\mathrm{sign}(H_{i})-0.5+0.5E_{\epsilon,i},$
@@
$A_{i}$ $\leftarrow X_{i},$
$Y_{i}$ $\leftarrow\beta(A_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
@@
$\beta(a)=1$ (homogeneous treatment effect),
$\beta(a)=2\exp(-a^{2}/2)$ (heterogeneous treatment effect),
@@
First, we simulate two datasets with $N=1000$ i.i.d. samples, one from setting (het.)/(Z lin.) and one from setting (het.)/(Z nonlin.) and apply the methods (het. T.E. mlIV) and (het. T.E. linearIV). We use two different bandwidths. The first one is the normal reference rule / Silverman’s rule of thumb [25, 34] and is given by $h_{N}=1.06\min(s,Q/1.34)N^{-1/5}$, where $s$ is the sample standard deviation and $Q$ the interquartile range of $\{A_{i}\}_{i=1}^{N}$. For valid inference, we need undersmoothing $h_{N}\ll N^{-1/5}$ (Assumption 4) and hence for the smaller bandwidth, we replace the $N^{-1/5}$ by $N^{-2/7}$ in the normal reference rule, which is also, what is done in [17]. The estimated heterogeneous treatment effect functions together with (pointwise) standard confidence intervals (dashed lines) and (pointwise) robust confidence sets (shaded region) can be seen in Figure 1. From the plots, we see that in the setting (Z lin.), the methods (het. T.E. mlIV) and (het. T.E. linearIV) yield almost identical results, both in terms of point estimate and confidence sets. In contrast, in the setting (Z nonlin.), the method (het. T.E. mlIV) seems to yield more stable results with smaller confidence bounds, whereas the results using (het. T.E. linearIV) are very unstable even in regions where there are a lot of samples $A_{i}$. This is consistent with our theory (see Remarks 6 and 11). Moreover, we also observe that the standard confidence intervals and the robust confidence sets are almost identical in regions with a lot of samples $A_{i}$ and large instrument strength. In contrast, if either there are not a lot of samples $A_{i}$ near a point $a$ (e.g. at the boundary) or the strength of the instrument is not very strong (e.g. method (het. T.E. linearIV) in the setting (Z nonlin.)), the robust confidence sets can be considerably larger than the standard confidence sets.
@@
We now want to investigate the efficiency of the different estimators and the coverage of the different confidence sets more extensively. For each combination of setting (hom.)/(het.) with (Z lin.)/(Z nonlin.) we consider sample size $N\in\{250,500,1000,1500,2000\}$ and simulate $n_{\text{sim}}=500$ datasets from (26). We still use generalized additive models (gam) for the nuisance functions and the Epanechnikov kernel for methods (het. T.E. mlIV) and (het. T.E. linearIV). Moreover, we only consider undersmoothing for the bandwidth, i.e. we use the normal reference rule times $N^{1/5}/N^{2/7}$. Let us first consider the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.). Here, we use the methods (hom. T.E. mlIV) and (hom. T.E. linearIV) as well as the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 2. As it is to be expected, the mean squared errors (MSEs) of the estimators decrease with $N$ and the MSEs of the (hom. T.E.) estimators are smaller than the MSEs of the comparable (het. T.E.) estimators. The MSEs of the (het. T.E.) estimators for $a=0$ are smaller than the MSEs for the (het. T.E.) estimators for $a=1.5$, which is due to the fact that the density of $A_{i}$ is larger at $a=0$ than at $a=1.5$. Moreover, we see that both the standard and robust confidence sets exhibit good empirical coverage close to the desired 0.95. Only the standard confidence intervals for (het. T.E.) at $a=1.5$ have a coverage that is a bit too low, again reflecting the lack of a sufficiently large number of samples near $a=1.5$. Notably, the robust confidence intervals also exhibit good coverage also in this regime. Regarding the difference between settings (Z lin.) and (Z nonlin.), we see that when the instrument $Z$ acts linearly on the treatment (Z lin.), the performance of linearIV and mlIV is very similar for both the homogeneous treatment effect and the heterogeneous treatment effect. On the other hand, we observe that in the setting (Z nonlin.), the methods based on mlIV have significantly lower MSE than their corresponding linearIV counterpart. This reflects the theoretical results that using the optimal machine learning instruments can lead to a significant efficiency gain.
@@
Figure 2: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $a=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $a=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=1.5$.
@@
Let us now consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.). Here, we only fit the methods that actually estimate a heterogeneous treatment effect. The results are shown in Figure 3. The overall picture is similar to before. Most notably, the coverage of the standard confidence interval for the (het. T.E.) methods at $a=1.5$ are a bit too low, but the coverage of the robust confidence sets is better. Moreover, in the setting (Z nonlin.), the methods based on mlIV have again lower MSEs than the methods based on linearIV.
@@
Figure 3: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $a=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $a=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $a=1.5$.
In Appendix E, we present further simulations where, we consider higher-dimensional covariates $X_{i}$.
@@ 4.3 Real Data Experiments
We apply our methods to two well-known instrumental variables datasets. Additionally to gam (generalized additive models using the R package mgcv [35, 36]), we also use boosted regression trees and random forest to estimate the nuisance functions. We use the R packages xgboost [9, 10] and ranger [38]. We refer to the three machine learning methods as gam, xgboost and random forest. In contrast to gam, for which we simply use the default settings, the methods xgboost and random forest need to be tuned. In view of the cross-fitting scheme, one would need to tune the machine learning methods for each of the $K$ different training datasets of size $N-N/K$ separately to avoid cross-contamination of the sample. However, recent research by [5] showed that this is usually not an issue and it suffices to tune the machine learning methods once on the full sample. Hence, in the following, for each nuisance function that has to be estimated, we tune the machine learner once on the full sample to obtain the optimal hyperparameters (with respect to cross-validated prediction accuracy for xgboost and with respect to out-of-bag prediction error for random forest) and then keep them fixed for the $K$ subsamples in all $S$ repetitions of the aggregation procedure from Section 4.1.
@@ 4.3.1 Effect of Institutions on Economic Growth
linearIV (xgboost) 0.62 0.35 $[-0.07,1.31]$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
@@
One might wonder if the homogeneous treatment effect is a realistic model assumption for the AJR dataset or if one should consider heterogeneity with respect to geography. We do not aim to give a comprehensive answer to this question, but for illustration purposes, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV), where we take the latitude as the variable $A_{i}$ with respect to which we want to quantify the heterogeneity. In Figure 4, we see the estimated heterogeneous effects with respect to latitude for the machine learning methods gam and xgboost. We omit the results for random forest as they look qualitatively similar. As for the initial example in Section 4.2, we use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing. From the plots, we cannot infer that there is a significant heterogeneity with respect to latitude. In fact, the robust confidence sets are unbounded for most values of latitude for both (linearIV) and (mlIV), especially when using xgboost for the nuisance functions and the smaller bandwidth. Note that due to the large difference between the robust confidence sets and the standard confidence intervals, we should not trust the standard confidence intervals. Especially, the size of the standard confidence intervals seems to decrease for large latitude. It is important to note, that this is an artifact of the data. For example, the confidence interval with latitude around $0.7$ is only based on the two datapoints with latitude close to $0.7$. Hence, there is no substantive value in the size of those confidence intervals. We conclude that for the AJR dataset, the sample size of $N=64$ is too small to make a sensible statement about a potential heterogeneous treatment effect.

### v2

@@ 3.3.1 Robust Confidence Sets
The following theorem ensures that under the null hypothesis, $\hat{Q}(\beta(v),v)$ is asymptotically normal.
@@
Under Assumptions 2, 3, 4, 5 and 6, we have that under $H_{0}:\beta(v)=\beta_{0}(v)$
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}\to\mathcal{N}(0,1)\text{ in distribution}.$
Moreover, the asymptotic variance of $\hat{Q}(\beta(v),v)$ can be consistently estimated, and hence, the approach for robust confidence sets is valid.
@@
Under the assumptions of Theorem 11 we have that under $H_{0}:\beta(v)=\beta_{0}(v)$
$\frac{\widehat{SE}_{Q}^{2}(\beta_{0}(v),v)}{\sigma_{Q}^{2}(v)}\to 1\text{ in probability.}$
@@
For the validity of the robust confidence sets, one still needs Assumption 5 which does not allow for IV that are too weak (see also Remark 3). However, we do not need the additional conditions $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(v),\,\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}^{2}(v)=o_{P}(h^{1/2})$ that – given a rate of convergence for $\hat{f}^{k}$ and $\hat{\phi}_{2}^{k}$ – pose even stronger restrictions on the IV-strength $\sigma_{Q}^{2}(v)$. In that sense, the robust confidence set construction works under slightly less restrictive conditions than the asymptotic normality of the point estimator.
@@ 4 Applications
We evaluate our method on simulated and real-world datasets. The method is implemented in the R package IVDML available on CRAN [38]. The code to reproduce the simulation and real data results is available on GitHub. [^1] https://github.com/cyrillsch/IVDML_Application Before considering simulations and real data applications, we point out an aspect that is important when applying the methods in practice.
@@ 4.1 Repeated Cross-Fitting
The partition of the dataset into the $K$ folds is random. The finite sample estimators $\hat{\beta}$ and $\hat{\sigma}^{2}$ for the homogeneous treatment effect and $\hat{\beta}(v)$ and $\hat{\sigma}^{2}(v)$ for the heterogeneous treatment effect depend on this random partition. Let us focus on the homogeneous treatment effect in the following (the aggregation for the heterogeneous treatment effect is analogous). Following [16], the procedure can be repeated $S$ times with independent random partitions, yielding estimators $(\beta_{s})_{s=1,\ldots,S}$ and $(\hat{\sigma}^{2}_{s})_{s=1,\ldots,S}$. The point estimator is then aggregated using the median, $\hat{\beta}^{*}=\mathrm{median}\{\hat{\beta}_{s};\,s=1,\ldots,S\}$ and the asymptotic variance additionally is updated by a term incorporating the sample variability, $\hat{\sigma}^{2,*}=\mathrm{median}\{\hat{\sigma}_{s}^{2}+(\hat{\theta}_{s}-\hat{\theta}^{*})^{2};\,s=1,\ldots,S\}$. For the robust confidence set construction, we use the same idea. Let $\left(\hat{Q}_{s}(\cdot)\right)_{s=1,\ldots,S}$ and $\left(\widehat{SE}_{Q,s}^{2}(\cdot)\right)_{b=1,\ldots,B}$ be obtained from (13) and (14). Define $\hat{Q}^{*}(\cdot)=\mathrm{median}\left\{\hat{Q}_{s}(\cdot);\,s=1,\ldots,S\right\}$ and $\widehat{SE}_{Q}^{2,s}(\cdot)=\mathrm{median}\left\{\widehat{SE}_{Q,s}^{2}(\cdot)+\left(\hat{Q}_{s}(\cdot)-\hat{Q}^{*}(\cdot)\right)^{2};\,s=1,\ldots,S\right\}$, where the medians are taken argument-wise.
@@ 4.2 Simulations
[…]
@@
$V_{i}$ $\leftarrow X_{i},$
$Y_{i}$ $\leftarrow\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$
@@
$\beta(v)=1$ (homogeneous treatment effect),
$\beta(v)=2\exp(-v^{2}/2)$ (heterogeneous treatment effect),
@@
First, we simulate two datasets with $N=1000$ i.i.d. samples, one from setting (het.)/(Z lin.) and one from setting (het.)/(Z nonlin.) and apply the methods (het. T.E. mlIV) and (het. T.E. linearIV). We use two different bandwidths. The first one is the normal reference rule / Silverman’s rule of thumb [40, 51] and is given by $h_{N}=1.06\min(s,Q/1.34)N^{-1/5}$, where $s$ is the sample standard deviation and $Q$ the interquartile range of $\{V_{i}\}_{i=1}^{N}$. For valid inference, we need undersmoothing $h_{N}\ll N^{-1/5}$ (Assumption 4) and hence for the smaller bandwidth, we replace the $N^{-1/5}$ by $N^{-2/7}$ in the normal reference rule, which is also, what is done in [24]. The estimated heterogeneous treatment effect functions together with (pointwise) standard confidence intervals (dashed lines) and (pointwise) robust confidence sets (shaded region) can be seen in Figure 1. From the plots, we see that in the setting (Z lin.), the methods (het. T.E. mlIV) and (het. T.E. linearIV) yield almost identical results, both in terms of point estimate and confidence sets. In contrast, in the setting (Z nonlin.), the method (het. T.E. mlIV) seems to yield more stable results with smaller confidence bounds, whereas the results using (het. T.E. linearIV) are very unstable even in regions where there are a lot of samples $V_{i}$. This is consistent with our theory (see Remarks 5 and 10). Moreover, we also observe that the standard confidence intervals and the robust confidence sets are almost identical in regions with a lot of samples $V_{i}$ and large instrument strength. In contrast, if either there are not a lot of samples $V_{i}$ near a point $v$ (e.g. at the boundary) or the strength of the instrument is not very strong (e.g. method (het. T.E. linearIV) in the setting (Z nonlin.)), the robust confidence sets can be considerably larger than the standard confidence sets.
@@
We now want to investigate the efficiency of the different estimators and the coverage of the different confidence sets more extensively. For each combination of setting (hom.)/(het.) with (Z lin.)/(Z nonlin.) we consider sample size $N\in\{250,500,1000,1500,2000\}$ and simulate $1000$ datasets from (26). We still use generalized additive models (gam) for the nuisance functions and the Epanechnikov kernel for methods (het. T.E. mlIV) and (het. T.E. linearIV). Moreover, we only consider undersmoothing for the bandwidth, i.e. we use the normal reference rule times $N^{1/5}/N^{2/7}$. Let us first consider the settings (hom.)/(Z lin.) and (hom.)/(Z nonlin.). Here, we use the methods (hom. T.E. mlIV) and (hom. T.E. linearIV) as well as the methods (het. T.E. mlIV) and (het. T.E. linearIV) for $\hat{\beta}(0)$ and $\hat{\beta}(1.5)$. The results are shown in Figure 2. As it is to be expected, the mean squared errors (MSEs) of the estimators decrease with $N$ and the MSEs of the (hom. T.E.) estimators are smaller than the MSEs of the comparable (het. T.E.) estimators. The MSEs of the (het. T.E.) estimators for $v=0$ are smaller than the MSEs for the (het. T.E.) estimators for $v=1.5$, which is due to the fact that the density of $V_{i}$ is larger at $v=0$ than at $v=1.5$. Moreover, we see that both the standard and robust confidence sets exhibit good empirical coverage close to the desired 0.95. Regarding the difference between settings (Z lin.) and (Z nonlin.), we see that when the instrument $Z$ acts linearly on the treatment (Z lin.), the performance of linearIV and mlIV is very similar for both the homogeneous treatment effect and the heterogeneous treatment effect. On the other hand, we observe that in the setting (Z nonlin.), the methods based on mlIV have significantly lower MSE than their corresponding linearIV counterpart. This reflects the theoretical results that using the optimal machine learning instruments can lead to a significant efficiency gain.
@@
Figure 2: Simulation results for setting (hom.)/(Z lin.) (top) and (hom.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}$ (hom. T.E.), $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Solid lines correspond to the methods (hom. T.E. linearIV) and (hom. T.E. mlIV). Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
@@
Let us now consider the settings (het.)/(Z lin.) and (het.)/(Z nonlin.). Here, we only fit the methods that actually estimate a heterogeneous treatment effect. The results are shown in Figure 3. The overall picture is similar to before. However, the coverage of the standard confidence interval for the (het. T.E.) methods at $v=1.5$ are a bit too low, but the coverage of the robust confidence sets is better. In the setting (Z nonlin.), the methods based on mlIV have again lower MSEs than the methods based on linearIV.
@@
Figure 3: Simulation results for setting (het.)/(Z lin.) (top) and (het.)/(Z nonlin.) (bottom). Left panel: mean squared error (MSE) of $\hat{\beta}(0)$ (het. T.E. for $v=0$) and $\hat{\beta}(1.5)$ (het. T.E. for $v=1.5$). Middle panel: coverage of standard confidence intervals. Right panel: coverage of robust confidence sets. Methods based on (linearIV) are in red and methods based on (mlIV) are in blue. Dashed lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=0$. Dotted lines correspond to the methods (het. T.E. linearIV) and (het. T.E. mlIV) for $v=1.5$.
In Appendix E, we present additional simulations: in Section E.1, we vary the IV strength and see that the standard confidence intervals may suffer from substantial undercoverage, whereas the robust confidence intervals still exhibit coverage close to the desired level of 0.95. In Section E.2, we consider higher-dimensional covariates and nuisance function estimation using boosted regression trees instead of generalized additive models.
@@ 4.3 Real Data Experiments
We apply our methods to two well-known instrumental variables datasets. Additionally to gam (generalized additive models using the R package mgcv [52, 53]), we also use boosted regression trees and random forest to estimate the nuisance functions. We use the R packages xgboost [12, 13] and ranger [55]. We refer to the three machine learning methods as gam, xgboost and random forest. In contrast to gam, for which we simply use the default settings, the methods xgboost and random forest need to be tuned. In view of the cross-fitting scheme, one would need to tune the machine learning methods for each of the $K$ different training datasets of size NN/KN-N/K separately to avoid cross-contamination of the sample. However, recent research by [5] showed that this is usually not an issue and it suffices to tune the machine learning methods once on the full sample. Hence, in the following, for each nuisance function that has to be estimated, we tune the machine learner once on the full sample to obtain the optimal hyperparameters (with respect to cross-validated prediction accuracy for xgboost and with respect to out-of-bag prediction error for random forest) and then keep them fixed for the $K$ subsamples in all $S$ repetitions of the aggregation procedure from Section 4.1.
@@ 4.3.1 Effect of Institutions on Economic Growth
linearIV (xgboost) 0.62 0.35 [0.07,1.31][-0.07,1.31] $(-\infty,\infty)$
@@
One might wonder if the homogeneous treatment effect is a realistic model assumption for the AJR dataset or if one should consider heterogeneity with respect to geography. We do not aim to give a comprehensive answer to this question, but for illustration purposes, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV), where we take the latitude as the variable $V_{i}$ with respect to which we want to quantify the heterogeneity. In Figure 4, we see the estimated heterogeneous effects with respect to latitude for the machine learning methods gam and xgboost. We omit the results for random forest as they look qualitatively similar. As for the initial example in Section 4.2, we use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing. From the plots, we cannot infer that there is a significant heterogeneity with respect to latitude. In fact, the robust confidence sets are unbounded for most values of latitude for both (linearIV) and (mlIV), especially when using xgboost for the nuisance functions and the smaller bandwidth. Note that due to the large difference between the robust confidence sets and the standard confidence intervals, we should not trust the standard confidence intervals. Especially, the size of the standard confidence intervals seems to decrease for large latitude. It is important to note, that this is an artifact of the data. For example, the confidence interval with latitude around $0.7$ is only based on the two datapoints with latitude close to $0.7$. Hence, there is no substantive value in the size of those confidence intervals. We conclude that for the AJR dataset, the sample size of $N=64$ is too small to make a sensible statement about a potential heterogeneous treatment effect.